OpenStudy (anonymous):

What is the distance between points (5,-3) and (0,9) in a coordinate system?

OpenStudy (anonymous):

its 13

OpenStudy (anonymous):

k how exactly do u do that

OpenStudy (anonymous):

x distance=5-0=5 y distance=-3-9=-12 according to pythagores theorem sqrt(5^2+((-12)^2)=13

OpenStudy (jiteshmeghwal9):

\[\LARGE{\sqrt{(a_2-a_1)^2+(b_2-b_1)^2}}\]

OpenStudy (anonymous):

thanks